OpenStudy (anonymous):

In a physics lab experiment, a spring clamped to the table shoots a 18g ball horizontally. When the spring is compressed 22cm , the ball travels horizontally 4.9m and lands on the floor 1.5m below the point at which it left the spring. What is the spring constant?

OpenStudy (anonymous):

does its velocity becomes zero after 4.9 m?

OpenStudy (anonymous):

frm mgh=1/2mv^2 .. find velocity... then acc. ... then f=ma=kx

OpenStudy (anonymous):

okay, found velocity to be 5.42 m/s

OpenStudy (anonymous):

how to find acceleration?

OpenStudy (souvik):

 the total mechanical energy is conserved... so \[mgh+1/2kx^2=1/2mv^2\]...............(1) where \(v\) is velocity when the ball hits the ground... now the initial horizontal velocity of the ball is zero so \(h=1/2gt^2\) or \(t=\sqrt{2h/g}\)......[ t= the time taken to fall] and \(v _{y}=g t\) and \(v _{x}=4.9/t\) now find v ...using..\[v=\sqrt{v _{x}^{2}+v _{y}^{2}}\] now put all the data you know in equation 1 and find the spring constant k....